Problem 5A-8 High-Low Method; Predicting Cost [LO5-10]

Nova Company’s total overhead cost at various levels of activity are presented below:

Month Machine-
Hours
Total
Overhead
Cost
April 46,000 $ 179,200
May 36,000 $ 156,700
June 56,000 $ 201,700
July 66,000 $ 224,200

Assume that the total overhead cost above consists of utilities, supervisory salaries, and maintenance. The breakdown of these costs at the 36,000 machine-hour level of activity is:

Utilities (variable) $ 43,200
Supervisory salaries (fixed) 61,000
Maintenance (mixed) 52,500
Total overhead cost $ 156,700

Nova Company’s management wants to break down the maintenance cost into its variable and fixed cost elements.

Required:

1. Estimate how much of the $224,200 of overhead cost in July was maintenance cost. (Hint: to do this, it may be helpful to first determine how much of the $224,200 consisted of utilities and supervisory salaries. Think about the behavior of variable and fixed costs.)

2. Using the high-low method, estimate a cost formula for maintenance in the form Y = a + bX.

3. Express the company’s total overhead cost in the form Y = a + bX.

4. What total overhead cost would you expect to be incurred at an activity level of 41,000 machine-hours?
Answer #1
1 Variable Utilities cost per Machine hour($43,200/36,000 hrs) $                                  1.20
Total Overhead Cost at 66,000 machine hours $                           2,24,200
Less:Utilities Costs(66,000*$1.20) $                              79,200
Less:Supervisory Salaries Costs $                              61,000
Maintaiance Costs $                              84,000
2 Maintainance Costs
Machine Hours Maintainance Costs
High level of activity                                  66,000 $84,000
Low level of activity                                  36,000 $52,500
Change                                  30,000 $31,500
Variable cost per MHrs =$31,500 / 30,000 =$1.05 per machine hrs
Fixed cost element =$84,000 - (66,000*$1.05)
Y =$14,700 + $1.05X
3 Total Overhead Costs
Machine Hours Overhead Costs
High level of activity                                  66,000 $2,24,200
Low level of activity                                  36,000 $1,56,700
Change                                  30,000 $67,500
Variable cost per MHrs =$67,500 / 30,000 =$2.25 per machine hrs
Fixed cost element =$224,200 - (66,000*$2.25)
Y =$75,700 + $2.25X
4 Cost to be incurred at 41,000 machine hrs =$75,700 + $2.25*41,000 =$167,950